OpenStudy (anonymous):

write an equation of the line parallel to y=8x-1 that contains (-6 2)

terenzreignz (terenzreignz):

The point-slope form. When two lines are parallel, they are to have the same slope, right? So what's the slope of y = 8x - 1 ?

OpenStudy (anonymous):

well I got (1,-7)

terenzreignz (terenzreignz):

That's a point... it's not even on the line y = 8x - 1.... Here's a hint If your line equation has the form y = mx + b then m is your slope So... what then is the slope of y = 8x - 1 ? ^.^

OpenStudy (anonymous):

sorry. my computer froze. Okay. So I got -8/1. Is that right?

terenzreignz (terenzreignz):

No... given your equation of the form y = mx + b m itself, is the slope... get it?

OpenStudy (anonymous):

oooooooh! Okay! I think I'm starting to understand a little bit.

terenzreignz (terenzreignz):

So... verdict? What's the slope of y = 8x - 1?

OpenStudy (anonymous):

is it 8? Sorry if I'm getting this wrong. I really am bad at math. I'm trying to double check my work.

OpenStudy (anonymous):

because 8/1=8?

terenzreignz (terenzreignz):

No, you're not bad at maths, just nervous, I reckon :) y = mx + b is something of a godsend when considering linear functions Here's an example... y = 3x + 4 3 is the slope y = -4x -9 -4 is the slope y = 2x - 218 2 is the slope getting it? :) When you have the form y = mx + b whatever is the coefficient (m) of x, THAT is the slope, regardless of b. So, can you now say with full confidence what the slope of y = 8x - 1 is? ^.^

OpenStudy (anonymous):

8 is the slope.

terenzreignz (terenzreignz):

That's correct. Now when two lines are parallel, they should have the same slope, right? So you're asked to construct a line, with slope 8, and passing through the point (-6 , 2) Any ideas how to go about this?

terenzreignz (terenzreignz):

(-6 , 2) sorry

OpenStudy (anonymous):

so would it be 8(-2/6)=

OpenStudy (anonymous):

I'm not too sure on how to solve this

OpenStudy (anonymous):

do i divide?

terenzreignz (terenzreignz):

Well, given a point (a,b) and a slope m the line having the slope m and passing through the point (a,b) has the equation y - b = m(x - a) OR y = mx - ma + b Now, just plug in :)

OpenStudy (anonymous):

y-1=8(6-2) is that it?

terenzreignz (terenzreignz):

How did you get y - 1? You have the point (-6 , 2) and the slope 8... so just plug in the formula y = mx - ma + b

OpenStudy (anonymous):

I think I've got this down now for my assignment. thanks so much for the help. :)

terenzreignz (terenzreignz):

No problem :)